OpenStudy (anonymous):

Can somebody please demonstrate what the substitution trick is for this? \[\huge\int\limits \cos(\sqrt{x}) dx\] (+1 medal for quality :-) )

OpenStudy (turingtest):

by parts I think :)

OpenStudy (unklerhaukus):

let \(u^2=x\)\[2u\cdot\text du=\text dx\] \[\Rightarrow\int\limits \cos u \cdot 2\text du\]

OpenStudy (unklerhaukus):

***\[\int\limits \cos(u)\cdot 2u \text du\]

OpenStudy (turingtest):

\[u=\sqrt x\]\[du=\frac12x^{-1/2}dx=\frac{dx}{2u}\implies dx=2udu\]gives\[\int2u\cos udu\]or am I wrong?

OpenStudy (turingtest):

guess not :)

OpenStudy (unklerhaukus):

the errors were all mine TT

OpenStudy (turingtest):

:P it happens

OpenStudy (unklerhaukus):

\(\int u\text dv=[uv]-\int v\text du\)

OpenStudy (anonymous):

^_^ (comes in to look at what's been going on and figure out) \[2 \int u\cos udu\] And obviously then u=u so du=du, so dv=cos(u) so v=sin(u)du I should be able to take it away from here thank you both, not sure whom to give the medal to though :D

OpenStudy (unklerhaukus):

du=1

OpenStudy (anonymous):

du = 1 du, so sure 1. :-)

OpenStudy (anonymous):

Kind of odd how the dx's or du's appear or disappear

OpenStudy (turingtest):

yeah, that's the right setup to be a little picky, I should point out that dv=cos(u)du du has to go somehwehere when we integrate ;) @UnkleRhaukus no, I think agentx is right \[u=u\implies du=du\]

OpenStudy (anonymous):

The "with respect to" part is always kind of quirky to me

OpenStudy (turingtest):

yes there is an invisible 1 in front, but who cares?

OpenStudy (turingtest):

yeah but it's not a full derivative; just a differential

OpenStudy (turingtest):

so there is no "with respect to" anything

OpenStudy (unklerhaukus):

i got confused , with the u's ., errors

OpenStudy (turingtest):

yeah, since we used a u-sub, using the variable u again to mean something else in the by parts formula is confusing we maybe should have started with a \(\theta\) substitution to avoid confusion :P

OpenStudy (turingtest):

...or some other letter

OpenStudy (anonymous):

That was a really clever trick back here when you had to find the derivative of \(\sqrt{x}\) and you get 2du = \(\sqrt{x}\) dx and then subsitute \(\sqrt{x}\) back in for u

OpenStudy (anonymous):

Thank you both for your time, I'm sure I'll be back. Getting these tips & tricks down is exactly what I was needing. Does no good to just know the answer

OpenStudy (turingtest):

a lot of integral that people think are unsolvable, or require a trig sub can be solved with a diligent u-sub as I did above do not underestimate the power of the u-sub if you follow it all the way through

OpenStudy (anonymous):

Got to know HOW or I'll be toast on the exam ^_^ (and not the good tasty toast, the burned smoldering toast)

OpenStudy (turingtest):

lol

OpenStudy (anonymous):

Turing I may have another along that line here in a moment, in fact I may have had one earlier that was solved via a different method. I'll PM you the link

OpenStudy (anonymous):

Err... I can't PM you @TuringTest lol

OpenStudy (turingtest):

okay :) I can try I will fan you first so you can message me

OpenStudy (turingtest):

now you can do it I think